Suppose that a gallon of wall paint costs $16 and covers 350 square feet. Three rooms of the house measure 9ft by 12 ft, 14ft by 16ft, and 12ft by 14 ft. The ceiling is 9ft high.
You wish to put a fawn colored carpet in the green room, a plum colored carpet in the off white room and a powder blue colored carpet in the blue room. Calculate the amount of each type of carpeting you would need. Suppose the carpeting is $18.50 per square yard. Calculate the cost of carpeting these rooms.You can't buy part of a yard so round your answer up to the next yard.

Answer by Abbey(339)

In the 9 x 12 room, there is a floor space of 9*12 feet or 3 x 4 yards (3 feet to a yard)
3*4 = 12 square yards
12 sq. yards * 18.50 per square yard = $222
Area is found by multiplying height by width:
Two walls are 9' high x 9' wide and two walls are 9'high x 12' wide
(2*9*9) + (2*9*12) = 162+216 = 378 square feet
350 square feet are in one gallon so you will need to purchase two gallons = $32.00 dollars
Total for room #1 = $254
Room 2
Same process
Walls: (2*9*14)+(2*9*16) = 540 sq. feet (2 gallons of paint = $32)
Floors: 14 * 16 = 224 square feet or 24.8 square yards (9 square feet in one yard) Round up to 25 square yards * 18.5 = $462.50
Total for room two = $494.50
Room 3
walls: (2*9*12)+(2*9*14) = 468 square feet - 2 gallons of paint = $32
carpet: (12*14=168 square feet or 18.67 square yards) Round up to 19 yards.
19*18.50 = $351.50
Total for room 3 = $383.50
Total for house = $383.50 + $462.50 + $254.00 = $1100.